OpenStudy (anonymous):

Integral of 1/(x+1)^2. This feels like it should be simple, but I can't put my finger on the answer...

OpenStudy (solomonzelman):

What is your best shot on the problem ?

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{w^n}~dw=\int\limits_{~}^{~}w^{-n}dw=\frac{1}{-n+1}w^{-n+1}+C}\)

OpenStudy (solomonzelman):

\(\large\color{black}{\displaystyle\int\limits_{~}^{~}\frac{1}{w^2}~dw=\int\limits_{~}^{~}w^{-2}dw=\frac{1}{-2+1}w^{-2+1}+C=-w^{-1}+C}\)

OpenStudy (anonymous):

I was thinking of product ruling it with (x+1)^-2 as u and dx as dv... but it doesn't seem like that'll go anywhere. The natural log antiderivat thing might work, but I haven't done it with a denominator to a power before...

OpenStudy (solomonzelman):

No, see what I wrote above? Do you agree with what I wrote?

OpenStudy (solomonzelman):

When (x+1) is your variable that makes no difference because dx would =du if you set u=x+1, and thus you can just treat x+1 as a single variable, or if you like you can perform that u sub anyway

OpenStudy (anonymous):

pellet, you're right. Goddamn substitution. Thanks.

OpenStudy (solomonzelman):

Anytime \(!\)